OpenStudy (anonymous):

A diamond merchant has a chance to buy one of 4 diamonds and wants to sell it immediately. One costs only $1000, but he has a 90% chance of getting $5000 for it (or not selling it at all). The second costs $10,000, and he has a 60% change of selling it for $25,000 (or not selling it at all). The third costs $20,000, but he has a 50% chance of selling it for $60,000 (or not selling it at all). The last diamond costs $50,000, but he has a 1% chance of selling it for $1,000,000 (or not selling it at all). Which diamond gives him the best expected value?

OpenStudy (anonymous):

options are : the one for $1000 the one for $10,000 the one for $20,000 the one for $50,000

OpenStudy (kropot72):

First convert the percentages to decimals, for example 90% converts to a probability of 0.9. Then multiply the profit by the probability to get the expected value. Cost Profit Probability Expected Value $1000 5000 - 1000 = $4000 0.9 $4000 * 0.9 = $3600 Now you can complete the table to find the correct choice of answer.

OpenStudy (kropot72):

@shavu1 Are you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait what chart

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

I have put the details for the first diamond in the above table to show you the method. Now you need to put in the details for the other three diamonds.

OpenStudy (anonymous):

am i looking for what expected value gives me 1mil

OpenStudy (anonymous):

1.3600 2.9000 3.20000 4.95000

OpenStudy (anonymous):

i calculated and this what i got for the last 3

OpenStudy (kropot72):



OpenStudy (anonymous):

i finished the chart

OpenStudy (anonymous):

wouldnt the right answer be D it has the highest expected value

OpenStudy (kropot72):

What was your calculated expected value for D?

OpenStudy (anonymous):

95,000

OpenStudy (kropot72):

Not correct. Check your calculation again.

OpenStudy (anonymous):

shouldnt it be 1,000,000-50,000= 95,000 x .01

OpenStudy (anonymous):

nvm D is 950

OpenStudy (anonymous):

the best answer would be C since its 20,000

OpenStudy (kropot72):

The calculation for D is : $950,000 * 0.01 = $9500

OpenStudy (anonymous):

so its C

OpenStudy (kropot72):

Yes, C gives the highest expected value.

OpenStudy (anonymous):

THANKS :d

OpenStudy (kropot72):

You're welcome :)